# EnvoDat: A Large-Scale Multisensory Dataset for Robotic Spatial Awareness and Semantic Reasoning in Heterogeneous Environments

### v1

@@ II Prior works
Several datasets have been introduced over the years [10], [11], [7], and each has its specific environmental focus and conditions. We reviewed the key attributes of these datasets, e.g., the environments in which they were collected, the sensory modalities they employed, and the specific conditions under which they were curated. We compared them against the EnvoDat dataset in Table I.
@@ II-A Urban, Indoor, and Natural Environment Datasets
The datasets presented in Table I are some of the most popular existing benchmarks. Majority of them focus on urban on-road environments, particularly for outdoor autonomous drive (e.g., KITTI [7], Oxford RobCar [8], CADC [18], etc.). Others focus on urban place recognition, e.g., MulRan [20], HeliPR [22], CODa [21], Cityscapes [15], Warburg et al. [23], etc. For indoor environments, datasets that capture long-term trajectories with loop closure and challenging factors like surface reflectivity, transparency, and dynamics [3], which are critical for sensors like LiDAR and RGB-D cameras are lagging. The TUM benchmarks [12], [9], ICL [24], VECtor [25], SUN RGB-D [14], and SceneNet RGB-D [26] often focus on small spatial (e.g., single room), fully observable, feature-dense environments with few or no dynamic objects. In Table I, we refer to these focus as spatial scale variation (SSV) and scene diversity (Div). High SSV and Div for indoor (I) and urban (U) environments include datasets with a wide range of spaces, from small rooms to large halls, and narrow alleys to wide streets. Low SSV and Div focus on uniform spaces, like only small rooms or similar-sized urban areas. For subterranean (S) and vegetated (V) environments, high SSV and Div cover varied spaces, such as narrow tunnels to large caverns, and different vegetation types from dense forests to open fields. Low SSV and Div cover similar-sized tunnels or specific vegetation types without much variation.
@@ III EnvoDat dataset - The overview
Sc.Typ NSeq LiDAR PtC LiDAR Img lay 2D Scan RGB+D ImS Mono CT Bag Dur NIm NAn OTC
@@
*Legends: Sc.Typ $\rightarrow$ Scene type, details of the individual scene, can be found at [29]. SubT $\rightarrow$ Underground tunnels. NSeq $\rightarrow$ Number of sequence. LiDAR PtC $\rightarrow$ 3D LiDAR point clouds data in pcd format at $10Hz$. LiDAR Img lay $\rightarrow$ Four 2D image data layers of the 3D LiDAR (the signal, reflectivity, range, and near-infra-red image frames) at $10Hz$. RGB+D $\rightarrow$ RGB and depth image frames of Intel Realsense D435i camera at $30fps$. ImS $\rightarrow$ IMU data samples at $100Hz$. Mono $\rightarrow$ RGB images of ELP 4K monocular camera at $30fps$. CT $\rightarrow$ Data collection time, M = morning, A = Afternoon, and E = Evening. Bag $\rightarrow$ Uncompressed ROS bag size in gigabytes (GB). Dur $\rightarrow$ Duration of raw data recording in hours. NIm $\rightarrow$ Number of annotated image frames. NAn $\rightarrow$ Number of annotations. OTC $\rightarrow$ Number of object and terrain classes.
@@ III-B Sensors Setup, Synchronisation and Calibration
In each scene of the EnvoDat, we captured raw sensor data using different sensor suites. The sensor specifications, setup and calibration are described in [29]. Each sensor in the suite transmits data at its own rate and with individual timestamps, which can cause slight timing variations. To address this, we employed the approximate time policy of the ROS [31] message filters framework to synchronise the timestamps of the incoming data across different sensory modalities. This approach ensured sub-millisecond accuracy and maintained time-synchronisation and the overall consistency of the fussed sensor data across all the scenes.
@@ III-D Ground-truth Poses
We provide high-precision reference ground-truth data for each scene, tailored to the scene’s features and geometric characteristics. Since most of the scenes in the EnvoDat such as the indoor and the sub tunnels are GPS denied, we obtained the pseudo-ground truth data from the fusion of different constraints derived from LiDAR inertial estimations from [33] and [34]. To verify the robustness and accuracy of these pseudo-ground truth trajectories, we compared the reconstructed LiDAR inertial odometry map with the ground truth geospatial map of the scene. Additionally, we assessed the scale precision by spatially reconstructing the map of the scene using the raw LiDAR sensor’s point cloud data. Figure 2(a-c) shows our quantitative and qualitative assessment, obtained by overlaying the global estimated trajectory poses on both the geospatial and the 3D spatial reconstructed maps.
@@ IV Experiments
  • RQ2 - To what extent do feature density or sparsity affect robotic autonomy and perception in heterogeneous environments?
@@ IV-A SLAM Algorithms Benchmark
To address RQ1, we benchmarked five SOTA SLAM algorithms - two visual-based (i.e., RTAB [36] and ORBSLAM3 [37]), two graph-based LiDAR SLAM (i.e., HDL [38] and GLIM [33]), and one filter-based LiDAR SLAM (i.e., FAST-LIO2 [34]) on our dataset. We selected five scenes from EnvoDat, which exhibit at least some of the geometric and feature characteristics highlighted in RQ1, (i): dynamic entities - MU-Hall-01, (ii): varying illumination and opaque surfaces - MU-TXN-01, (iii): zero and partial visibility conditions - SubT-ZAB-01 and MU-Cor-03 (night sequence), and (iv): high-dimensional observations & features-sparsity - Leo-Str-01. We analysed the impact of these environmental factors on the mapping accuracy by comparing the mapped points and estimated trajectories to the ground truth data (see Table III). We defined per-point absolute trajectory error (ATE), relative pose error (RPE) and scale drift (SD) as evaluation criteria.
@@
Table III shows the benchmark results of the SLAM algorithms, with the best metrics in bold. The majority of the algorithms struggled with the complexities of the environments, resulting in high ATE and varying SD values. GLIM [33] performed nearly optimally in all the scenes, regardless of the environments’ features and geometric characteristics. However, the visual-based methods failed in the subterranean and night sequences due to poor visibility conditions.
@@
Method Metrics Dynamic Entities V.Illum. & Opaque Zero & Partial Visibility HD Observ. & Sparsity
@@
GLIM [33] ATE ($\mu/\sigma$) 0.19 / 0.12 1.11 / 0.74 0.50 / 3.49 0.18 / 0.16 0.16 / 0.10
RPE ($\mu/\sigma$) 0.01 / 0.00 0.00 / 0.00 0.17 / 0.00 0.00 / 0.00 0.02 / 0.00
SD ($\mu/\sigma$) 0.99 / 0.04 0.99 / 0.05 0.99 / 0.06 1.00 / 0.02 0.98 / 0.12
@@
$\mu\rightarrow$ mean, $\sigma\rightarrow$ standard deviation. V.Illum. $\rightarrow$ Varying illumination. HD Observ. & Sparsity $\rightarrow$ High dimensional observations and feature sparsity.
@@ IV-B Feature Sparsity and Density
To establish the correlation between the feature density and the errors (ATE & RPE), first, we map the trajectory points to feature density bins. For each trajectory point, e.g., $\hat{\mathbf{t}}_{i}=(\hat{x}_{i},\hat{y}_{i},\hat{z}_{i})$, we find the bin index $(x_{\text{bin},i},y_{\text{bin},i},z_{\text{bin},i})$ for each coordinate with the corresponding edges $(\hat{x}_{i}\text{ in }\mathbf{x}_{\text{edges}},\hat{y}_{i}\text{ in }\mathbf{y}_{\text{edges}},\hat{z}_{i}\text{ in }\mathbf{z}_{\text{edges}})$. Thereafter, we assign the feature density at the trajectory point as: $\mathbf{d}_{i}=\rho(\textbf{x}_{\text{bin},i},\textbf{y}_{\text{bin},i},\textbf{z}_{\text{bin},i})~{}\text{if bin indices are valid}$, and 0 otherwise. $\mathbf{d}=[d_{1},d_{2},\ldots,d_{n}]$ is the feature density at the trajectory points, $\rho(.)$ is the feature density of the voxel that contains the trajectory point. For simplicity, we assume that out-of-bounds points have zero features. Finally, we compute the 2D Pearson correlation between the errors and the feature densities $\mathbf{d}$ as $\rho_{\text{ATE},\mathbf{d}}=\text{corr}(\text{ppATE},\mathbf{d})$ and $\rho_{\text{RPE},\mathbf{d}}=\text{corr}(\text{ppRPE},\mathbf{d})$.
@@
IV-C Object Detection and Classification Benchmarks
We address RQ3 using the EnvoDat dataset by evaluating three pre-trained object detector models: YOLOv8 [39], Fast R-CNN [40], and Detectron2 [41]. We trained these models on the annotated RGB images drawn across all the scenes in the EnvoDat dataset, which exhibit the characteristics outlined in RQ3. For consistency, we trained all the models with equal hyperparameter settings (e.g., learning rate $=0.00025$, batch size $=16$, epochs $=150$). We evaluated their performances based on accuracy, precision, and efficiency.
[…]
@@ IV-C Object Detection and Classification Benchmarks
For precision, we utilised the precision and recall metrics, similar to [42] and [22]. To balance these two metrics, we also computed the F1-score, the harmonic mean of precision and recall. Further, for efficiency, we evaluated the models in terms of inference time ($t_{\text{inf}}$), inference speed (frame per second i.e., $\text{FPS}=1/t_{\text{inf}}$), and GPU memory usage. Table IV shows the performance of these models on EnvoDat. The environment heterogeneity had a noticeable impact on their results. Nevertheless, YOLOv8 outperformed other models across most metrics, particularly in mAP and inference speed. Detectron2 and Fast R-CNN lagged significantly, especially in inference time and memory usage. Thus, YOLOv8 can be considered a more efficient choice for real-time applications.
@@ V Conclusion
We introduced EnvoDat, a large-scale multisensory dataset for advancing robots’ autonomy, spatial awareness and semantic reasoning in heterogeneous environments. EnvoDat covers a range of indoor, outdoor, and subterranean environments with varying geometric and feature characteristics. We benchmarked several SOTA SLAM and supervised learning algorithms on EnvoDat, showing how real-world conditions affect their performance in challenging environments, including multi-degraded, dynamic, and feature-sparse areas. We plan to expand our dataset by capturing more sequences, adding new scenes and providing more fine-grained layered polygon-based annotations to accelerate the development of real-time ecocentric perception and SLAM algorithms.
ACKNOWLEDGMENT
This project has received funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - No #430054590 (TRAIN).

### v2

@@ II Prior works
Several datasets have been introduced over the years [10, 11, 7], and each has its specific environmental focus and conditions. We reviewed the key attributes of these datasets, e.g., the environments in which they were collected, the sensory modalities they employed, and the specific conditions under which they were curated. We compared them against the EnvoDat dataset in Table I.
@@ II-A Urban, Indoor, and Natural Environment Datasets
The datasets presented in Table I are some of the most popular existing benchmarks. The majority of them focus on urban on-road environments, particularly for the outdoor autonomous drive (e.g., KITTI [7], Oxford RobCar [8], CADC [18], etc.). Others focus on urban place recognition, e.g., MulRan [20], HeliPR [22], CODa [21], Cityscapes [15], Warburg et al. [23], etc. For indoor environments, datasets that capture long-term trajectories with loop closure and challenging factors like surface reflectivity, transparency, and dynamics [3], which are critical for sensors like LiDAR and RGB-D cameras are lagging. The TUM benchmarks [12], [9], ICL [24], VECtor [25], SUN RGB-D [14], and SceneNet RGB-D [26] often focus on small spatial (e.g., single room), fully observable, feature-dense environments with few or no dynamic objects. In Table I, we refer to these focus as spatial scale variation (SSV) and scene diversity (Div). High SSV and Div for indoor (I) and urban (U) environments include datasets with a wide range of spaces, from small rooms to large halls, and narrow alleys to wide streets. Low SSV and Div focus on uniform spaces, like only small rooms or similar-sized urban areas. For subterranean (S) and vegetated (V) environments, high SSV and Div cover varied spaces, such as narrow tunnels to large caverns, and different vegetation types from dense forests to open fields. Low SSV and Div cover similar-sized tunnels or specific vegetation types without much variation.
@@ III EnvoDat dataset - The overview
Sc.Typ NSeq LiDAR PtC LiDAR Img lay 2D Scan RGB+D ImS Mono CT Size Dur NIm NAn OTC
@@
*Legends: Sc.Typ $\rightarrow$ Scene type, details of the individual scene, can be found at [29]. SubT $\rightarrow$ Underground tunnels. NSeq $\rightarrow$ Number of sequence. LiDAR PtC $\rightarrow$ 3D LiDAR point clouds data in pcd format at $10Hz$. LiDAR Img lay $\rightarrow$ Four 2D image data layers of the 3D LiDAR (the signal, reflectivity, range, and near-infra-red image frames) at $10Hz$. RGB+D $\rightarrow$ RGB and depth image frames of Intel Realsense D435i camera at $30fps$. ImS $\rightarrow$ IMU data samples at $100Hz$. Mono $\rightarrow$ RGB images of ELP 4K monocular camera at $30fps$. CT $\rightarrow$ Data collection time, M = morning, A = Afternoon, and E = Evening. Size $\rightarrow$ Estimated download size including the uncompressed ROS bag files in gigabytes (GB). Dur $\rightarrow$ Duration of raw data recording in hours. NIm $\rightarrow$ Number of annotated image frames. NAn $\rightarrow$ Number of annotations. OTC $\rightarrow$ Number of object and terrain classes.
@@ III-B Sensors Setup, Synchronisation and Calibration
In each scene of the EnvoDat, we captured raw sensor data using different sensor suites. The sensor specifications, setup and calibration are described in [29]. Each sensor in the suite transmits data at its own rate and with individual timestamps, which can cause slight timing variations. To address this, we employed the approximate time policy of the ROS [31] message filters framework to synchronise the timestamps of the incoming data across different sensory modalities. This approach ensured a few seconds of accuracy and maintained time-synchronisation and the overall consistency of the fussed sensor data across all the scenes.
@@ III-D Ground-truth Poses
We provide high-precision reference ground-truth data for each scene, tailored to the scene’s features and geometric characteristics. Since most of the scenes in the EnvoDat such as the indoor and the sub tunnels are GPS denied, we obtained the pseudo-ground truth data from the fusion of different constraints derived from LiDAR inertial estimations from GLIM [33]. To verify the robustness and accuracy of these pseudo-ground truth trajectories, we compared the reconstructed LiDAR inertial odometry map with the ground truth geospatial map of the scene. Additionally, we assessed the scale precision by spatially reconstructing the map of the scene using the raw LiDAR sensor’s point cloud data. Figure 2(a-c) shows our quantitative and qualitative assessment, obtained by overlaying the global estimated trajectory poses on both the geospatial and the 3D spatial reconstructed maps.
@@ IV Experiments
  • RQ2 - To what extent does the feature density or sparsity influence robotic autonomy and scene perception in heterogeneous environments?
@@ IV-A SLAM Algorithms Benchmark
To address RQ1, we benchmarked five SOTA SLAM algorithms on our dataset. We selected two visual-based methods (RTAB [35] and ORB-SLAM3 [36]), one graph-based LiDAR method (HDL-SLAM [37]), and two filter-based LiDAR methods (FAST-LIO2 [38] and DLIO [39]). We chose five scenes from EnvoDat, which exhibits at least some of the geometric and feature characteristics highlighted in RQ1: (i) dynamic entities - MU-Hall-01, (ii) varying illumination and opaque surfaces - MU-TXN-01, (iii) zero and partial visibility conditions - SubT-ZAB-01 and MU-Cor-03 (night sequence), and (iv) high-dimensional observations & features-sparsity - Leo-Str-01. We analysed the impact of these environmental factors on the mapping accuracy by comparing the mapped points and estimated trajectories to the ground truth data (see Table III). We defined per-point absolute trajectory error (ATE), relative pose error (RPE) and scale drift (SD) as evaluation criteria.
@@
Table III shows the benchmark results of the SLAM algorithms, with the best metrics in bold. Most of the algorithms struggled with the complexities of the environments. However, regardless of the environments’ features and geometric characteristics, Fast-LIO2 [38] and DLIO [39] consistently outperformed the other algorithms across all the scenes, with comparatively lower ATE and RPE values. Conversely, the visual-based methods failed in the subterranean and night sequences due to poor visibility conditions. The HDL-SLAM [37] also showed competitive results in some of the scenes; however, it does not match the overall robustness of Fast-LIO2 [38], which stands out as the top-performing algorithm in the benchmark.
@@
Methods Metrics Dynamic Entities V.Illum. & Opaque Zero & Partial Visibility HD Obs. & F-Spar
@@
DLIO [39] ATE ($\mu/\sigma$) 1.14 / 0.49 2.64 / 1.14 6.93 / 8.81 3.29 / 2.47 1.76 / 1.65
RPE ($\mu/\sigma$) 0.04 / 0.00 0.03 / 0.00 0.29 / 0.00 0.04 / 0.00 0.05 / 0.00
SD ($\mu/\sigma$) 1.10 / 0.50 2.19 / 6.34 1.45 / 3.94 1.19 / 3.19 1.12 / 2.05
@@
$\mu\rightarrow$ mean, $\sigma\rightarrow$ standard deviation. V.Illum. $\rightarrow$ Varying illumination. HD Obs. & F-Spar $\rightarrow$ High dimensional observations and feature sparsity.
@@ IV-B Feature Sparsity and Density
To establish the correlation between the feature density and the errors (ATE & RPE), first, we map the trajectory points to feature density bins. For each point, e.g., $\hat{\mathbf{t}}_{i}=(\hat{x}_{i},\hat{y}_{i},\hat{z}_{i})$, we find the bin index $(x_{\text{bin},i},y_{\text{bin},i},z_{\text{bin},i})$ for each coordinate with the corresponding edges $(\hat{x}_{i}\text{ in }\mathbf{x}_{\text{edges}},\hat{y}_{i}\text{ in }\mathbf{y}_{\text{edges}},\hat{z}_{i}\text{ in }\mathbf{z}_{\text{edges}})$. Thereafter, we assign the feature density at the point as: $\mathbf{d}_{i}=\rho(\textbf{x}_{\text{bin},i},\textbf{y}_{\text{bin},i},\textbf{z}_{\text{bin},i})~{}\text{if bin indices are valid}$, and 0 otherwise. $\mathbf{d}=[d_{1},d_{2},\ldots,d_{n}]$ is the feature density at the trajectory points, $\rho(.)$ is the feature density of the voxel that contains the point. For simplicity, we assume that out-of-bounds points have zero features. Finally, we compute the 2D Pearson correlation between the errors and the feature densities $\mathbf{d}$ as $\rho_{\text{ATE},\mathbf{d}}=\text{corr}(\text{ppATE},\mathbf{d})$ and $\rho_{\text{RPE},\mathbf{d}}=\text{corr}(\text{ppRPE},\mathbf{d})$.
@@
IV-C Supervised Learning Algorithms Benchmark
We address RQ3 using the EnvoDat by evaluating three pre-trained object detector models: YOLOv8 [40], Fast R-CNN [41], and Detectron2 [42]. We trained these models on the annotated RGB images drawn across all the scenes in the EnvoDat, which exhibit the characteristics outlined in RQ3. For consistency, we trained all the models with equal hyperparameter settings (e.g., learning rate $=0.00025$, batch size $=16$, epochs $=150$), with a 70% train, 20% validation, and 10% test split ratios. We evaluated their performances based on accuracy, precision, and efficiency.
For accuracy, we used the mean average precision (mAP) metric at multiple intersections over union (IoU) thresholds, $0.5$ and $0.5-0.95$. Formally, $\text{mAP}_{0.5}=\frac{1}{N}\sum_{i=1}^{N}\text{AP}_{i}^{(\text{IoU}=0.5)}$ and $\text{mAP}_{(0.5:0.95)}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{10}\sum_{t=0.5}^{0.95}\text{AP}_{i}^{(\text{IoU}=t)}$, where $N$ is the total number of object classes, and $\text{AP}_{i}$ is the average precision for class $i$.
@@ IV-C Supervised Learning Algorithms Benchmark
For precision, we utilised the precision and recall metrics, similar to [43] and [22]. To balance these two metrics, we also computed the F1-score, the harmonic mean of precision and recall. For efficiency, we evaluated the models in terms of inference time ($t_{\text{inf}}$), speed (frame per second i.e., $\text{FPS}=1/t_{\text{inf}}$), and GPU memory usage.
Table IV shows the models’ performance on EnvoDat. The environment heterogeneity had a noticeable impact on their results. Nevertheless, YOLOv8 outperformed other models across most metrics, particularly in mAP and inference speed. Detectron2 and Fast R-CNN lagged, especially in inference time and memory usage. Thus, YOLOv8 can be considered a more efficient choice for real-time applications.
@@ V Conclusion
We introduced EnvoDat, a large-scale multimodal dataset for advancing robotic autonomy, spatial awareness and semantic reasoning in heterogeneous environments. EnvoDat covers a range of indoor, outdoor, and subterranean environments with distinct geometric and feature characteristics. We benchmarked several SOTA SLAM and supervised learning algorithms on EnvoDat, showing how real-world conditions influence their performance in challenging environments, including multi-degraded, dynamic, and feature-sparse areas. We plan to expand our dataset by capturing more sequences, adding new scenes, and providing more fine-grained layered polygon-based annotations to accelerate the development of real-time ecocentric perception and SLAM algorithms.